a chain of lenght L is lying on a table with 1/3 of its lenght hanging down. Calculate the work done in pulling the chain up to the surface of table

Grade:9

1 Answers

vikas askiitian expert
13 years ago

this question can easily be solved if we consider the motion of center of mass of hanging part....

at intial moment L/3 length was hanging , so the distance of center of mass of this hanging part the from table is L/6...

mass of this hanging part = M/3

finally when the chain is pulled up then center of mass of hanging part is on table ....

workdone = mgdh                        (dh = change in height)

m = M/3 , h = L/6

work = M/3 *g* L/6

      =MGL/18

.